OpenStudy (anonymous):

Solve for t. Write your answer to 3 decimal places. 2500 = 500 e^.06t. Help! Will give medal!!!!

OpenStudy (anonymous):

can someone help me on this?

OpenStudy (ikram002p):

ok now devide both side on 500 2500=500 e^0.06t 500=e^0.06t take ln for both side ln(500)=ln(e^0.06t) ln500=0.06 t now solve for t t=(ln 500)/0.06

OpenStudy (anonymous):

t = 103.5768 or t = 103.577?

OpenStudy (anonymous):

nvm thank you!! :)